OpenStudy (anonymous):

Part A: Max rented a motorbike at $465 for 5 days. If he rents the same motorbike for a week, he has to pay a total rent of $625. Write an equation in the standard form to represent the total rent (y) that Max has to pay for renting the motorbike for x days.

OpenStudy (anonymous):

@aaronq @Loser66

OpenStudy (mathstudent55):

$625 for a week $465 for 5 days A week is how many days?

OpenStudy (anonymous):

7

OpenStudy (mathstudent55):

Good, so we can rewrite what we have as: $625 for 7 days $465 for 5 days Now we subtract the 2 lines to get: $160 for 2 days Ok so far?

OpenStudy (anonymous):

OK

OpenStudy (mathstudent55):

If it is $160 for 2 days, then dividing by 2, we get $80 per day

OpenStudy (mathstudent55):

If it costs $80 per day, how much is the charge for 5 days? What is 5 * $80 = ?

OpenStudy (anonymous):

400

OpenStudy (mathstudent55):

Great, but the charge of r 5 days was $465, not $400. That means there is a fixed fee plus a daily fee. We know the daily fee is $80. The fixed fee must be the excess of $65 from $400 to $465.

OpenStudy (mathstudent55):

That means the equation is: y = 80x + 65

OpenStudy (mathstudent55):

Now we can check. For x = 5 (5 days), y = 80 * 5 + 65 = 465 For x = 7 (7 days, a week), y = 80 * 7 + 65 = 625 Both rental periods give us the correct cost according to the given info, so this is the correct equation: y = 80x + 65

OpenStudy (anonymous):

Thanks